# Revitalizing Multivariate Time Series Forecasting: Learnable Decomposition with Inter-Series Dependencies and Intra-Series Variations Modeling

### v2

@@ 4.3 Model Analysis
We observe substantial performance improvements in Table 3 introduced by the ‘Auto-regressive self-attention’ and ‘Channel-wise self-attention’ components. ‘Auto-regressive self-attention’ brings an average of 19.02% of MSE to decrease across five datasets compared to using linear layer, and ‘Channel-wise self-attention’ achieves 21.09% improvement. Moreover, their synergistic integration yields further enhancements in model performance, getting an average 25.03% of MSE decrease, reaching an optimal level. This attests to the efficacy of both design elements and once again validates our hypothesis, namely, appropriately modelling inter-series dependencies and intra-series variations in multivariate time series can yield better predictive performance.
@@
Superiority of trainable 1D convolutional kernel over moving average kernel. To better emphasize the advantages of our trainable 1D convolutional kernel with weights initialized using a Gaussian distribution, in comparison to conventional moving average kernels, we conducted an extensive experiment. Given that DLinear arguably represents the most prominent instantiation utilizing a moving average kernel for trend information extraction, and achieves performance comparable to other state-of-the-art methods with the mere use of two simple linear layers, we select it as our baseline. For comparison, we substitute its moving average kernel with our trainable 1D convolutional kernel, denoting the modified model as LD_TL if the kernel is set to trainable, else LD_UTL if the kernel is set to untrainable.
@@
In comparison to a simple moving average kernel, the trainable 1D convolutional kernel, as depicted in Table 4, consistently exhibits superior predictive performance across all four datasets regardless of its trainability. The untrainable version of 1D convolutional kernel brings an average 7.28% of MSE decrease across five datasets compared to using a moving average kernel, while the trainable version gives 11.98%. The obtained results conclusively demonstrate the superior efficacy of our trainable 1D convolutional kernel over a simple moving average kernel, and the trainability of the kernel plays a significant role in its adaptability. We leave further analysis to APPENDIX B.1-B.3.
@@
Analysis of different attention mechanisms in inter-series dependencies modelling. We devised a comprehensive experiment to investigate three prevalent approaches for modelling inter-series dependencies: ‘Channel-wise self-attention’, ‘Point-wise self-attention’, and ‘Patch-wise self-attention’. The first considers the entire sequence of a variate as a token (Liu et al., 2024), the second regards distinct variables at the same timestamp as tokens (Zhou et al., 2022a, c), and the third treats patches of the raw sequence as tokens (Zhang & Yan, 2023; Nie et al., 2023). Specifically, we eliminate the ‘Auto-regressive self-attention’ branch from the original Leddam structure to concentrate on inter-series dependencies modelling. We sequentially employ ‘Channel-wise self-attention’, ‘Point-wise self-attention’, and ‘Patch-wise self-attention’. In Figure 4, it is readily apparent that compared to ‘Point-wise self-attention’, and ‘Patch-wise self-attention’, ‘Channel-wise self-attention’ exhibits superior predictive performance, implies a much better inter-series dependencies modelling ability.
@@
Analysis of different attention mechanisms in intra-series variations modelling. Similarly, we replace the ‘Channel-wise self-attention’ branch from the original Leddam structure and use ‘Point-wise self-attention’, ‘Patch-wise self-attention’, and ‘Auto-regressive self-attention’ for intra-series variations modelling. The superiority of the ‘Auto-regressive self-attention’ architecture is clearly proved by the experimental results in Figure 4.
@@ 4.4 Learnable Decomposition Generalization Analysis
Model selection and experimental setting. To achieve this objective, we conduct experiments across a spectrum of representative time series forecasting model structures, including (1) Transformer-based methods: Informer (Zhou et al., 2022a), Transformer (Vaswani et al., 2017); (2) Linear-based methods: LightTS (Zhang et al., 2022); and (3) TCN-based methods: SCINet (LIU et al., 2022); (4) RNN-based methods: LSTM (Hochreiter & Schmidhuber, 1997). We standardize the input length $T$ to 96, and similarly, the prediction length $F$ is uniformly set to 96. Subsequently, comparative experiments were conducted on five datasets: ETTh2, ETTm2, Weather, Electricity, and Traffic. Specifically, we sequentially replaced the ‘Auto-regressive self-attention’ and ‘Channel-wise self-attention’ components in Leddam with each model. The comparative analysis was performed to assess the predictive performance before and after the incorporation of Leddam, comparing the original models with the augmented counterparts. And more results can be found in APPENDIX C.
Quantitative results. In Table 5, it is apparent that the incorporation of the Leddam structure leads to a notably substantial enhancement in the predictive performance of various models, even with the introduction of only a single linear layer. Specifically, LightTS demonstrates an average MSE reduction of 11.87% across five datasets, other models are LSTM: 48.56%, SCINet: 23.15%, Informer: 31.72%, and Transformer: 26.27%. Particularly noteworthy is the performance enhancement observed in the classical LSTM model, where the MSE experiences a remarkable decrease of 76.97% and 80.28% on ETTh2 and ETTm2, respectively, a result that is profoundly surprising. This unequivocally substantiates the generality of the Leddam structure.
@@ A.2 IMPLEMENTATION DETAILS AND MODEL PARAMETERS
We trained our Leddam model using the L2 loss function and employed the ADAM optimizer. We initialized the random seed as $rs=2021$. We also configured the hyperparameter $K=25$-kernel size of the decomposition kernel (the Moving Average Kernel (MOV) and learnable decomposition module (LD)). During the training process, we incorporated an early stopping mechanism, which would halt training after six epochs if no significant reduction in loss was observed on the validation set. For evaluation purposes, we used two key performance metrics: the mean square error (MSE) and the mean absolute error (MAE). We carried a grid hyperparameter search, where dimension of layer $D\in\{256,512\}$, learning_rate $lr\ in\{0.001,0.0001,0.0005\}$, dropout ratio $dr\in\{0.0,0.2,0.5\}$ and number of network layers $nl\in\{1,2,3\}$. Our implementation was carried out in PyTorch and executed on an NVIDIA V100 32GB GPU. All the code and experimental details will be public in the future. We also attach the code in a supplementary file for your reference and run the demo. All the compared baseline models that we reproduced are implemented based on the benchmark of TimesNet (Wu et al., 2023) Repository, which is fairly built on the configurations provided by each model’s original paper or official code.
@@ Appendix B FURTHER ANALYSIS OF DIFFERENT DECOMPOSITION METHODOLOGIES
Figure 5: Visualization of different decomposition schemes in ETTh1, ETTh2, ETTm1, and ETTm2 data. RAW means the raw time series. MOV means Moving average kernel, LD means our learnable decomposition module.
Figure 6: Visualization of different decomposition schemes in Electricity, Solar-Energy, Traffic, and Weather data. RAW means the raw time series. MOV means Moving average kernel, LD means our learnable decomposition module.
Figure 7: Visualization of decomposition results before (Untrained) and after (Trained) training of learnable decomposition module (LD) in Electricity, Solar-Energy, Traffic, and Weather data.
B.1 DECOMPOSITION RESULTS VISUALIZATION
We present a detailed comparison of results obtained by decomposing various datasets using the Moving Average Kernel (MOV) and learnable decomposition module (LD) in Figure 56. The decomposition performance of LD is consistently superior to that of the MOV across all eight datasets.
We also present a detailed comparison of results obtained by decomposing various datasets using the learnable decomposition module (LD) before and after training in Figure 7. It’s evident that our LD has demonstrated enhanced adaptability in decomposing time series after training.
@@ B.2 VISUALIZATION OF WEIGHTS
We present visualizations of the weight for the Learnable Decomposition Module (LD) and Moving Average Kernel (MOV). It is observed that, unlike MOV, which employs uniform weights across all datasets, our LD adapts to the characteristics of different datasets, generating context-specific weights for sequence decomposition.
B.3 FURTHER IMPROVEMENT OF MULTI-SCALE HYBRID DECOMPOSITION
$i$ $=1,\ldots,n$
$X_{Trend}^{i}$ $=AvgPool(Padding(X_{embed}))_{kernel_{i}}$
$X_{Trend}$ $=Intergrating(X_{Trend}^{1},\ldots,X_{Trend}^{n})$
$X_{Seasonal}$ $=X_{embed}-X_{Trend}$ (8)
We also conducted exploratory experiments to investigate the performance of our Learnable Decomposition Module (LD) in this particular context. We replaced the fundamental decomposition kernel employed in MICN, i.e., the basic Moving Average Kernel (MOV), with our LD. Subsequently, comparative experiments were conducted on four datasets(ETTm2, Weather, Electricity, Traffic) with input length $T=96$ and prediction length $F=96$. The LD consistently outperform MOV under a multi-scale decomposition strategy across all datasets in Table 6, which highlights the excellence of LD.
Appendix C GENERALIZATION OF Leddam
To better illustrate the generality of our Leddam framework and its performance improvement across various models, we opted to visualize the predictive outcomes on three representative datasets (Electricity, ETTh2, Traffic) in Figure 9-13. For a prediction task with input length $T=96$ and prediction length $F=96$, we present a comparative analysis of the predictive performance of different models before and after integrating our Leddam framework. To achieve this objective, we conduct experiments across a spectrum of representative time series forecasting model structures, including (1) Transformer-based methods: Informer (Zhou et al., 2022a), Transformer (Vaswani et al., 2017); (2) Linear-based methods: LightTS (Zhang et al., 2022); and (3) TCN-based methods: SCINet (LIU et al., 2022); (4) RNN-based methods: LSTM (Hochreiter & Schmidhuber, 1997). After the introduction of the Leddam framework, various models have consistently demonstrated superior predictive performance.
@@ Appendix C GENERALIZATION OF Leddam
Appendix D INFLUENCE OF INPUT LENGTH ON PREDICTION PERFORMANCE
Figure 14: Forecasting performance (MSE and MAE) of Leddam with varying look-back windows on 4 datasets: ETTm2, Electricity, Traffic, and Weather. The look-back windows are selected to be $T\in\{48,96,192,336,504,720,960\}$, and the prediction horizons are $F=720$.
Appendix E FULL FORECASTING RESULTS

### v5

@@ 4.3 Model Analysis
We observe substantial performance improvements in Table 3 introduced by the ‘Auto-regressive self-attention’ and ‘Channel-wise self-attention’ components. ‘Auto-regressive self-attention’ brings an average of 19.02% of MSE to decrease across five datasets compared to using linear layer, and ‘Channel-wise self-attention’ achieves 21.09% improvement. Moreover, their synergistic integration yields further enhancements in model performance, getting an average 25.03% of MSE decrease, reaching an optimal level. This attests to the efficacy of both design elements and once again validates our hypothesis, namely, appropriately modeling inter-series dependencies and intra-series variations in multivariate time series can yield better predictive performance.
@@
Superiority of Learnable Decomposition Module over Moving Average Kernel. To better emphasize the advantages of our Learnable Decomposition Module with weights initialized using a Gaussian distribution, in comparison to conventional Moving Average Kernel, we conducted an extensive experiment. Given that DLinear arguably represents the most prominent instantiation utilizing a Moving Average Kernel for trend information extraction, and achieves performance comparable to other state-of-the-art methods with the mere use of two simple linear layers, we select it as our baseline. For comparison, we substitute its Moving Average Kernel with our Learnable Decomposition Module, denoting the modified model as LD_TL if the kernel is set to trainable, else LD_UTL if the kernel is set to untrainable.
@@
In comparison to a simple Moving Average Kernel, the Learnable Decomposition Module, as depicted in Table 4, consistently exhibits superior predictive performance across all four datasets regardless of its trainability. The untrainable version of the 1D convolutional kernel brings an average 7.28% of MSE decrease across five datasets compared to using a Moving Average Kernel, while the trainable version gives 11.98%. The obtained results conclusively demonstrate the superior efficacy of our Learnable Decomposition Module over a simple Moving Average Kernel, and the trainability of the kernel plays a significant role in its adaptability. We leave further analysis to APPENDIX B.1-B.4.
Decomposition result analysis. To further investigate why our Learnable Decomposition Module (LD) is a better time series decomposition solution than Moving Average Kernel (MOV), we conducted the following analysis on the seasonal part and trend part obtained.
Since the seasonal part represents repetitive patterns in the raw sequence, a good seasonal part should capture all major frequencies in the raw sequence. We separately calculated the amplitude similarity of the dominant frequencies (top 25%) between the seasonal part obtained by each method and the raw sequence, denoted as FFT. A better decomposition strategy should yield a seasonal part with higher similarity to the dominant frequencies of the raw sequence.
A good trend part should effectively capture the trend changes in the original sequence. Thus, we employed Dynamic Time Warping (DTW) (Mü, 2007) to compute the similarity between the raw sequence and the trend parts obtained from the two decompositions, denoted as DTW. A superior decomposition strategy should result in a trend part with higher DTW similarity to the raw sequence.
We selected the final variate of each dataset. The variates used are as follows.
  • ETT Dataset: Oil Temperature (OT) every hour or 15 minutes.
  • Electricity: hourly electricity consumption of the 321st (last) user.
  • Solar-Energy: solar power production every 10 minutes of the 137th (last) PV plant.
  • Traffic: hourly road occupancy rates measured by the 862nd (last) sensor.
  • Weather: CO2 (ppm) collected every 10 minutes.
LD is pre-trained on task: input-96-forecast-720. To avoid cherry-picking, both metrics are computed by averaging the results calculated over the entire test dataset.
Table 5: Decomposition result analysis. Comparison of DTW and FFT for LD and MOV across eight datasets.
Kernel LD (Ours) MOV
Dataset/Metric DTW FFT DTW FFT
Electricity 0.643 0.942 0.618 0.931
Solar_Energy 0.910 0.781 0.873 0.734
Traffic 0.603 0.993 0.563 0.991
Weather 0.858 0.760 0.846 0.691
ETTh1 0.741 0.892 0.724 0.852
ETTh2 0.675 0.900 0.652 0.887
ETTm1 0.821 0.754 0.808 0.717
ETTm2 0.925 0.894 0.908 0.759
The DTW and FFT of LD are consistently superior to that of the MOV across all eight datasets in Table 5. This demonstrates that LD is a better time series decomposition method compared to MOV.
@@
Analysis of different attention mechanisms in inter-series dependencies modeling. We devised a comprehensive experiment to investigate three prevalent approaches for modeling inter-series dependencies: ‘Channel-wise self-attention’, ‘Point-wise self-attention’, and ‘Patch-wise self-attention’. The first considers the entire sequence of a variate as a token (Liu et al., 2024), the second regards distinct variables at the same timestamp as tokens (Zhou et al., 2022a, c), and the third treats patches of the raw sequence as tokens (Zhang & Yan, 2023; Nie et al., 2023). Specifically, we eliminate the ‘Auto-regressive self-attention’ branch from the original Leddam structure to concentrate on inter-series dependencies modeling. We sequentially employ ‘Channel-wise self-attention’, ‘Point-wise self-attention’, and ‘Patch-wise self-attention’. In Figure 4, it is readily apparent that compared to ‘Point-wise self-attention’, and ‘Patch-wise self-attention’, ‘Channel-wise self-attention’ exhibits superior predictive performance, implies a much better inter-series dependencies modeling ability.
@@
Analysis of different attention mechanisms in intra-series variations modeling. Similarly, we replace the ‘Channel-wise self-attention’ branch from the original Leddam structure and use ‘Point-wise self-attention’, ‘Patch-wise self-attention’, and ‘Auto-regressive self-attention’ for intra-series variations modeling. The superiority of the ‘Auto-regressive self-attention’ architecture is proved by the experimental results in Figure 4.
@@ 4.4 Learnable Decomposition Generalization Analysis
Model selection and experimental setting. To achieve this objective, we conduct experiments across a spectrum of representative time series forecasting model structures, including (1) Transformer-based methods: Informer (Zhou et al., 2022a), Transformer (Vaswani et al., 2017); (2) Linear-based methods: LightTS (Zhang et al., 2022); and (3) TCN-based methods: SCINet (LIU et al., 2022); (4) RNN-based methods: LSTM (Hochreiter & Schmidhuber, 1997). We standardize the input length $T$ to 96, and similarly, the prediction length $F$ is uniformly set to 96. Subsequently, comparative experiments were conducted on five datasets: ETTh2, ETTm2, Weather, Electricity, and Traffic. Specifically, we sequentially replaced the ‘Auto-regressive self-attention’ and ‘Channel-wise self-attention’ components in Leddam with each model. The comparative analysis was performed to assess the predictive performance before and after the incorporation of Leddam, comparing the original models with the augmented counterparts. And more results can be found in APPENDIX D.
Quantitative results. In Table 6, it is apparent that the incorporation of the Leddam structure leads to a notably substantial enhancement in the predictive performance of various models, even with the introduction of only a single linear layer. Specifically, LightTS demonstrates an average MSE reduction of 11.87% across five datasets, other models are LSTM: 48.56%, SCINet: 23.15%, Informer: 31.72%, and Transformer: 26.27%. Particularly noteworthy is the performance enhancement observed in the classical LSTM model, where the MSE experiences a remarkable decrease of 76.97% and 80.28% on ETTh2 and ETTm2, respectively, a profoundly surprising result. This unequivocally substantiates the generality of the Leddam structure.
@@ 5 Conclusion
Acknowledgements
This work was partially supported by the Start-up Fund of The Hong Kong Polytechnic University (No. P0045999) and the Seed Fund of the Research Institute for Smart Ageing (No. P0050946). We thank anonymous reviewers for their comments and suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.2 Implementation Details and Model Parameters
We trained our Leddam model using the L2 loss function and employed the ADAM optimizer. We initialized the random seed as $rs=2021$. We also configured the hyperparameter $k=25$-kernel size of the decomposition kernel (the Moving Average Kernel (MOV) and Learnable Decomposition Module (LD)). During the training process, we incorporated an early stopping mechanism, which would halt training after six epochs if no significant reduction in loss was observed on the validation set. For evaluation purposes, we used two key performance metrics: the mean square error (MSE) and the mean absolute error (MAE). We carried a grid hyperparameter search, where dimension of layer $dim\in\{256,512\}$, learning rate $lr\ in\{0.001,0.0001,0.0005\}$, dropout ratio $dr\in\{0.0,0.2,0.5\}$ and number of network layers $nl\in\{1,2,3\}$. Our implementation was carried out in PyTorch and executed on an NVIDIA V100 32GB GPU. All the compared baseline models that we reproduced are implemented based on the benchmark of TimesNet (Wu et al., 2023) Repository, which is fairly built on the configurations provided by each model’s original paper or official code. It is worth noting that both the baselines used in this paper and our Leddam have fixed a long-standing bug. This bug was originally identified in Informer (Zhou et al., 2022a) (AAAI 2021 Best Paper) and subsequently addressed by FITS (Anonymous, 2024a). For specific details about the bug and its resolution, please refer to GitHub Repository [^4] https://github.com/VEWOXIC/FITS.
@@ Appendix B Further Analysis of Different Decomposition Methodologies
B.1 Decomposition Result Visualization
Figure 5: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on ETTh1.
Figure 6: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on ETTh2.
Figure 7: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on ETTm1.
Figure 8: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on ETTm2.
Figure 9: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on Electricity.
Figure 10: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on Solar-Energy.
Figure 11: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on Traffic.
Figure 12: Trend-Seasonal Decomposition Results obtained by LD (Red) and MOV (Blue) on Weather.
We present a detailed comparison of results obtained by decomposing various datasets using the Moving Average Kernel (MOV) and Learnable Decomposition Module (LD).
We still selected the final variate of each dataset. LD are pretrained on task: input-96-forecast-720. Given a sampling frequency of 10, 15 minutes, or 1 hour for these datasets, we opt to decompose time series of two lengths: 120 and 720, to better contrast the decomposition results and reflect the extracted seasonal and trend patterns respectively.
The decomposition performance of LD is consistently superior to that of the MOV across all eight datasets in Figure 512.
@@ B.2 Visualization of Weights
We present visualizations of the weight for the Learnable Decomposition Module (LD) and Moving Average Kernel (MOV) in Figure 13. It is observed that, unlike MOV, which employs uniform weights across all datasets, our LD adapts to the characteristics of different datasets, generating context-specific weights for sequence decomposition.
B.3 Impact of RevIN
To explore how much ReVIN (Kim et al., 2022) improves the performance of LD (and MOV), we utilized Autoformer (Wu et al., 2021) as the baseline and investigated the performance of LD and MOV under scenarios with and without ReVIN. In Table 7, We computed that ReVIN yielded an average MSE performance improvement (reduction) of 5.24% across all tasks for LD, and 9.04% for MAE. For MOV: 1.75% for MSE and 3.32% for MAE. Compared to MOV, ReVIN would bring greater performance improvements to LD.
Table 7: Model Comparison
Dataset/Metric pred_len LD+ReVIN MOV+ReVIN LD MOV
MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.667 0.526 0.694 0.538 0.664 0.531 0.671 0.534
720 0.659 0.538 0.753 0.559 0.647 0.547 0.688 0.556
ETTm2 96 0.207 0.281 0.218 0.292 0.216 0.299 0.223 0.306
720 0.414 0.404 0.422 0.41 0.435 0.43 0.426 0.423
Weather 96 0.201 0.249 0.217 0.266 0.289 0.366 0.233 0.304
720 0.373 0.369 0.378 0.373 0.402 0.408 0.394 0.396
B.4 Further Improvement of Multi-scale Hybrid Decomposition
$i$ $=1,\ldots,n$
$X_{Trend}^{i}$ $=AvgPool(Padding(X_{embed}))_{kernel_{i}}$
$X_{Trend}$ $=Intergrating(X_{Trend}^{1},\ldots,X_{Trend}^{n})$
$X_{Seasonal}$ $=X_{embed}-X_{Trend}$ (8)
We also conducted exploratory experiments to investigate the performance of our Learnable Decomposition Module (LD) in this particular context. We replaced the fundamental decomposition kernel employed in MICN, i.e., the basic Moving Average Kernel (MOV), with our LD. Subsequently, comparative experiments were conducted on four datasets(ETTm2, Weather, Electricity, Traffic) with input length $T=96$ and prediction length $F=96$. The LD consistently outperforms MOV under a multi-scale decomposition strategy across all datasets in Table 8, which highlights the excellence of LD.
@@ B.4 Further Improvement of Multi-scale Hybrid Decomposition
Appendix C Further Analysis of Auto-regressive Self-attention
Our experiments in Table 9 demonstrate that our Auto-regressive Embedding, compared to Point-wise and Patch-wise Embedding, maintains crucial temporal positional information within permutation-invariant self-attention mechanisms. This is demonstrated by the low dependency of Auto-regressive Embedding on positional encoding.
Table 9: Forecasting performance of Auto-regressive, Point-wise, and Patch-wise Embedding on the ETTh2, ETTm2, and Weather datasets. The forecasting task is input-96-forecast-96/720. w/o means without position embedding, pos means position embedding is used.
Models pred_len Type Auto Patch Point
Dataset/Metric MSE MAE MSE MAE MSE MAE
ETTh2 96 w/o 0.294 0.343 0.382 0.429 1.453 1.000
pos 0.293 0.343 0.354 0.404 0.958 0.803
720 w/o 0.413 0.432 0.926 0.703 3.533 1.625
pos 0.416 0.433 0.835 0.666 1.569 0.991
ETTm2 96 w/o 0.176 0.257 0.226 0.321 0.419 0.511
pos 0.175 0.256 0.208 0.314 0.335 0.447
720 w/o 0.398 0.395 0.963 0.726 2.826 1.386
pos 0.393 0.394 0.732 0.671 1.411 0.996
Weather 96 w/o 0.166 0.211 0.176 0.240 0.219 0.309
pos 0.167 0.213 0.165 0.234 0.174 0.260
720 w/o 0.342 0.343 0.355 0.405 0.351 0.395
pos 0.342 0.344 0.337 0.370 0.346 0.388
Appendix D Generalization Analysis of Leddam
To better illustrate the generality of our Leddam framework and its performance improvement across various models, we opted to visualize the predictive outcomes on three representative datasets (Electricity, ETTh2, Traffic) in Figure 14-18. For a prediction task with input length $T=96$ and prediction length $F=96$, we present a comparative analysis of the predictive performance of different models before and after integrating our Leddam framework. To achieve this objective, we conduct experiments across a spectrum of representative time series forecasting model structures, including (1) Transformer-based methods: Informer (Zhou et al., 2022a), Transformer (Vaswani et al., 2017); (2) Linear-based methods: LightTS (Zhang et al., 2022); and (3) TCN-based methods: SCINet (LIU et al., 2022); (4) RNN-based methods: LSTM (Hochreiter & Schmidhuber, 1997). After the introduction of the Leddam framework, various models have consistently demonstrated superior predictive performance.
@@ Appendix D Generalization Analysis of Leddam
Appendix E Model Efficiency Study
We evaluated the parameter count, and the inference time (average of 5 runs on a single NVIDIA V100 32GB GPU) with $batch\_size=1$ on ETTh1 and Electricity dataset. We set the dimension of layer $dim\in\{96,192,336,720\}$, and the number of network layers $nl=2$. The task is input-96-forecast-720. We explored Leddam and four cutting-edge transformer-based multivariate time series forecasting models:iTransformer, Crossformer, PatchTST, and FEDformer. Results can be found in Table 10.
Table 10: Model efficiency analysis. * means ‘former.’ Para means ‘Parameter count(M).’ Time means ‘inference time(ms).’
Datasets/Models dim Leddam PatchTST Cross* iTrans* FED*
Param Time Para Time Para Time Para Time Para Time
ETTh1 256 2.50 233.92 3.27 251.00 8.19 399.00 1.27 177.67 3.43 303.556
512 9.20 249.34 8.64 266.66 32.11 445.74 4.63 190.92 13.68 345.736
Electricity 256 2.59 283.04 3.27 322.53 13.66 432.40 1.27 192.12 4.24 347.634
512 9.36 296.70 8.64 411.96 43.04 507.54 4.63 249.60 15.29 398.599
Appendix F Hyperparameter Sensitivity Analysis
We investigated the impact of the most significant parameters of Leddam: dimension of layer ($dim$), number of network layers ($nl$), dropout ratio ($dr$), and size of decomposition kernel ($k$). The default settings were: $dr=0.0$, $dim=512$, $nl=2$, $k=25$. Based on Table 11– 14, we can easily conclude that Leddam is insensitive to dropout ratio and kernel_size, while more sensitive to the number of network layers and the dimension of layers. For the kernel size, we argue after our initialization with a Gaussian distribution of 00 mean and $1$ variance, the weight of the central element is the largest. As moving away from the center, the weights gradually decrease. Therefore, when the kernel size increases, the weights of the farther points become smaller, thus not significantly affecting the final decomposition result.
Table 11: Impact of dropout ratio ($dr$). $dr\in\{0.0,0.1,0.2,0.5\}$
Dataset ETTm1 ETTm2
$dr$/Metric MSE MAE MSE MAE
0.0 0.469 0.447 0.406 0.400
0.1 0.468 0.446 0.404 0.401
0.2 0.476 0.452 0.406 0.401
0.5 0.474 0.452 0.409 0.404
Table 12: Impact of dimension of layer ($dim$). $dim\in\{128,256,512,1024\}$
Dataset ETTm1 ETTm2
$dim$/Metric MSE MAE MSE MAE
128 0.486 0.454 0.399 0.397
256 0.473 0.449 0.403 0.399
512 0.469 0.447 0.406 0.400
1024 0.492 0.459 0.402 0.397
Table 13: Impact of number of network layers ($nl$). $nl\in\{1,2,3,4\}$
Dataset ETTm1 ETTm2
$nl$/Metric MSE MAE MSE MAE
1 0.474 0.447 0.398 0.398
2 0.469 0.447 0.406 0.400
3 0.482 0.456 0.427 0.413
4 0.485 0.456 0.407 0.401
Table 14: Impact of kernel_size ($k$). $k\in\{15,25,55,75,105\}$
Dataset ETTm1 ETTm2
$k$/Metric MSE MAE MSE MAE
15 0.469 0.447 0.406 0.400
25 0.469 0.447 0.406 0.400
55 0.469 0.448 0.407 0.401
75 0.470 0.446 0.406 0.401
105 0.469 0.447 0.406 0.400
Appendix G Full Forecasting Results